The random variable X represents the number of heads that occur when you toss 4 coins at the same time what is the probability distribution



1
Expert's answer
2021-02-24T07:58:09-0500

Four coins are tossed. Therefore sample space contain "2^4" i.e "16" points.

Let "S" be the sample space.

Then "S=" { "HHHH,HHHT,HHTH,HHTT,HTHH,HTHT,HTTH,HTTT"

"THHH,THHT,THTH,THTT,TTHH,TTHT,TTTH,TTTT" }

Let "X" be the random variable representing the number of heads that occur.

Then "X" can take the values "0,1,2,3,4."

Now corresponding to each points of random variable "Z" we will get their probabilities.

"\therefore P(X=0)=" ​probability of getting no head

"=\frac{1}{16}"

"P(X=1)=" probability of getting one head

"=\frac{4}{16}=\frac{1}{4}"

"P(X=2)=" probability of getting two head

"=\frac{6}{16}=\frac{3}{8}"

"P(X=3)=" probability of getting three head

"=\frac{4}{16}=\frac{1}{4}"

"P(X=4)=" probability of getting four head

"=\frac{1}{16}"

Let "F(x)" be the distribution function of the random variable "X" .

If "-\infty<x<0," "F(x)=0"

If "0\leq x<1," "F(x)=P(X=0)=\frac{1}{16}" If "1\leq x<2, F(x)=[P(X=0)+P(X=1)]=(\frac{1}{16}+\frac{4}{16})=\frac{5}{16}"

If "2\leq x<3, F(x)=[P(X=0)+P(X=1)+P(X=2)]=(\frac{1}{16}+\frac{4}{16}+\frac{6}{16})=\frac{11}{16}"

If "3\leq x<4, F(x)=[P(X=0)+P(X=1)+P(X=2)+P(X=3)]=(\frac{1}{16}+\frac{4}{16}+\frac{6}{16}+\frac{4}{16})=\frac{15}{16}"

If "4\leq x<\infty , F(x)=[P(X=0)+P(X=1)+P(X=2)+P(X=3)+P(X=4)]"

"=(\frac{1}{16}+\frac{4}{16}+\frac{6}{16}+\frac{4}{16}+\frac{1}{16})=1"

Which is the required distribution function.